OpenStudy (anonymous):

Determine whether the geometric series is converges or diverges, if convergent find the sum of the series ∑_(n=1)^∞▒〖4(〖1/2)〗^k 〗

OpenStudy (experimentx):

use equation editor http://www.codecogs.com/latex/eqneditor.php

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

it converges r<1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what is the sum?

OpenStudy (anonymous):

please

OpenStudy (jhannybean):

\[\large \sum_{n=1}^{\infty}4(\frac12)^k\] does n=k? or is k supposed to be n? O_o

OpenStudy (experimentx):

at least write your equation well ... I can't say anything, i don't understand what you mean .. is that the above .. your question?

OpenStudy (jhannybean):

Use the p-series test p= 1/2 \[\large p \le 1 \implies D\]\[\large p > 1\implies C\]

OpenStudy (anonymous):

thank you

OpenStudy (jhannybean):

Np.

OpenStudy (anonymous):

this is the questions

OpenStudy (jhannybean):

Can you please post the problem? :) Not sure i want to save your attachments on my computer.

OpenStudy (anonymous):

can you solve it